### v3

# GlucOS: Security, correctness, and simplicity for automated insulin delivery

Abstract.
We present GlucOS, a novel system for trustworthy automated insulin delivery. Fundamentally, this paper is about a system we designed, implemented, and deployed on real humans and the lessons learned from our experiences.
GlucOS combines algorithmic security, driver security, and end-to-end verification to protect against malicious ML models, vulnerable pump drivers, and drastic changes in human physiology. We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy. Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize. Our results show that GlucOS maintains safety and improves glucose control even under attack conditions. This work demonstrates the potential for secure, personalized, automated healthcare systems. Our source code is open source.
1. Introduction
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1). However, insulin is a dangerous hormone as too much insulin can kill people in a matter of hours (cryer2012severe, ) and too little insulin can kill people in a matter of days (benoit2018trends, ). For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels. In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors. Several commercial (beta_bionics, ; tandem_iq, ; omnipod, ) and open-source (loop, ; openaps, ) automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
(a) Devices for automated insulin delivery. This figure is from breakthrought1d.org.
(b) Closed-loop architecture.
(c) Overall GlucOS architecture.
Figure 1. Overview of automated insulin delivery devices, closed-loop architecture, and GlucOS.
Current open-source automated insulin delivery systems are complex. For example, the OpenAPS (openaps, ) core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements. All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging. Automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, adding to the overall complexity and increasing the attack surface. In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code. None of the open-source automated insulin delivery systems uses formal methods, can withstand attacks on components, or has mechanisms or policies for security.
In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS. Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure ML-driven insulin delivery, (3) the ability to withstand attacks on insulin pump driver software, and (4) the application of formal methods to prove correct our insulin delivery path, the most sensitive part of the system.
Despite these efforts, we identify scenarios in automated insulin delivery that require users to take action for security. Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions.
Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery. We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
Automated insulin delivery systems are different from traditional applications that the systems and security communities focus on, but they are important. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design. Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few (cve1, ; cve2, ; cve3, ; cve4, ). However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people. Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
Our main contribution lies in a new end-to-end system that we built and the lessons learned from our experience designing, implementing, and deploying GlucOS to humans to manage their T1D. Our novel contributions include:
  • Security mechanisms and policies that protect individuals from malicious ML, vulnerable pump drivers, and drastic changes in the human’s glycemic response.
  • A case study describing our experiences from a real-world deployment with seven people, and results from simulation to show that our techniques generalize.
  • A human-centric design that considers the user as an integral part of the system, influencing both the architecture and including them in our defensive strategy.
We give top priority to the health, safety, and ethics of the humans using GlucOS to manage their T1D. Section 10.1 summarizes our ethical considerations.
There are 8.4 million people living with T1D worldwide (t1d_numbers, ), underscoring the need for providing trustworthy systems for automated insulin delivery. Our cross-disciplinary work, carried out with rigorous medical and ethical oversight, is a first step in addressing this need.
2. Overview
This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery; we have two primary goals. First, we want the software to be simple and correct. Second, we want to identify the most vulnerable parts of the system and design a system to prevent attacks on these components or withstand successful attacks. This section describes our overall architecture for GlucOS that strives to achieve these goals, and our threat model.
2.1. Locking down the insulin delivery path
We focus our system architecture on providing security mechanisms and policies on the insulin delivery path. The insulin delivery path includes models for calculating the amount of insulin still active in the individual (insulin can take up to six hours to fully absorb), closed-loop algorithms for calculating insulin dosing, deep neural network ML to predict future metabolic states, and drivers for programming the insulin pump. Together, these components are both the most complex, and the most sensitive from a security perspective because they are responsible for insulin delivery.
2.2. Architecture
From a system design perspective, we strive to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts. For our software system architecture, we use separation principles from the OS and microkernel areas (accetta1986mach, ; hartig1997performance, ) applied to the application layer to push complexity away from the most critical parts of the system. This architecture is similar to secure web browsers (grier2008secure, ; wang2009multi, ; tang2010trust, ; reis2019site, ), which also apply OS separation principles at the application layer. Figure 1(c) shows how we decompose our automated insulin delivery system, where we strive to keep our trusted computing base simple. GlucOS includes a BioKernel that forms our core trusted computing base. The BioKernel logs events that other components can ingest to learn the state of the system. These other components include apps for training ML models, visualizing statistics, an AI agent for suggesting dosing during meals, and an app for replaying closed-loop algorithm execution. Together, these components make up the overall GlucOS system.
Figure 2. Overall architecture for the BioKernel.
Figure 2 shows our architecture for the BioKernel. The BioKernel is the component that interacts with the continuous glucose monitor (CGM) and insulin pump hardware, executes the closed-loop dosing algorithm, runs security and safety checks, and produces event logs that other components use to learn the state of the system.
The BioKernel runs the closed-loop algorithm every five minutes as the new CGM readings come in. Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive ML model. The predictive ML model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states. The BioKernel then issues commands to the pump to deliver the calculated amount of insulin. A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
To withstand attacks, the BioKernel serves as a reference monitor (anderson1972computer, ) to enforce security policies on the predictive ML model’s pump commands and the pump driver’s low-level I/O. This reference monitor architecture ensures that we can withstand attacks from malicious predictive ML as well as malicious pump drivers without having to know the internals of how they operate.
For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a simple reactive safe model. For the pump driver, we interpose on the I/O path to ensure that only the pump commands issued by the BioKernel are programmed into the pump. These checks constitute our pump invariants, which, when violated switch the pump to operate in a safe mode.
Our use of formal methods focuses on the software related to delivering insulin, which is the most risky operation in any automated insulin delivery system. We use Hoare logic to prove key functions correct, define system states and transitions to handle runtime verification check failures safely, and we introduce the notion of a novel biological invariant. Our biological invariant is an end-to-end check of our theoretical calculations of how the individual’s metabolism should behave compared to what we observe in practice, where violations of this invariant suggest a potentially catastrophic failure of the system.
2.3. Threat model
The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia. Hypoglycemia results from an overdose of insulin. Once an automated insulin delivery system has delivered insulin, there is no way to remove it from the body. If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours. Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system. Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% (AmericanDiabetesAssociation2021, ).
As our design philosophy centers around humans, providing flexibility for people to use whatever ML they want by loading personalized models or using any pump hardware they want is important. This flexibility helps guide our focus on providing novel security mechanisms and policies for these components. Our goal is to make sure that they stay safe even if the model they pick is malicious or the pump driver they choose has security vulnerabilities. Attackers may target automated insulin delivery systems because they can directly harm the human using the system.
We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity. Thus, we do not consider attacks from tampered CGM sensors or pump hardware in our threat model. These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
We consider attacks from malicious predictive ML algorithms. These attacks can come from trojan attack logic embedded directly in a model, or from an adversarial attack on predictive ML. For adversarial attacks, although we assume that the CGM and insulin pump remain attack free, an adversary can manipulate these entities by serving an individual who is using our system with calculated amounts of carbohydrates in their food (e.g., at a restaurant), thus influencing the resulting CGM readings and corresponding insulin commands. In our threat model, malicious predictive ML can issue arbitrary insulin doses.
We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O to add, drop, or modify commands sent to the pump.
Successful attacks on predictive ML or pump drivers, if unchecked, can be devastating because they directly affect insulin delivery and can be used by an attacker to induce potentially life-threatening hypoglycemia.
3. Algorithmic security: Withstanding malicious ML
Currently, no automated insulin delivery systems use deep neural network ML. Some current systems use statistical methods, like linear regression or model predictive control (physio2011, ; physio2019, ; lunze2013control, ; control2019, ; openaps, ). The reason for using this more traditional form of ML is sound. It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions. However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ). Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” (chatgpt_hallucination, ) could have lethal consequences if it delivers an inappropriate insulin dose (cryer2012severe, ; benoit2018trends, ).
In this section, we discuss the design and implementation of GlucOS’s algorithmic security mechanisms to enable predictive models, including deep neural network ML, to run safely in automated insulin delivery systems.
3.1. Mechanism fundamentals
Our key novel insight underpinning our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a sufficiently long period. People inject insulin to facilitate glucose absorption after eating. Digestion converts food to glucose, and insulin facilitates the absorption of glucose to tissue (e.g., the brain) where the body can use it as an energy source. Since insulin dosing is driven by food consumption, all correct algorithms should dose the same amount of insulin to cover glucose from digestion.
However, the timing of this insulin delivery matters for maintaining balanced glucose levels (Section 9). Insulin activation is a slow process, with injected insulin taking up to six hours to fully activate. To meet this challenge, our algorithmic security mechanism pairs two models. Our first model is a predictive model that anticipates future metabolic states and doses insulin proactively. Our second model is a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently. This pairing allows us to balance proactive insulin dosing for faster responses with a more conservative approach for safety.
Architecturally, in a novel approach for automated insulin delivery systems, the predictive model runs outside of our trusted computing base while the reactive safe model is part of the trusted computing base. This separation enables individuals to use whatever predictive models they want, while maintaining security. It also defends against malicious predictive models, as all insulin delivery decisions are ultimately vetted by the reactive safe model. This design also enables people to update their predictive models without violating core safety guarentees.
Conceptually, the predictive model is the primary model in our system and should control the pump most of the time. We use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline. Then, we bound the size of this deviation. By combining these two models, we strive to get beneficial properties from both. Our approach combines the safety of a reactive safe model and the performance from a predictive model.
Our security mechanism draws inspiration from previous work on combining simple and complex models for safety (sha2001using, ; mohan2013s3a, ). However, our approach extends these ideas by providing a novel security mechanism specifically designed for insulin delivery models. Our key contributions lie in our insight into the security properties of insulin delivery algorithms and our principles for how to apply our mechanism to automated insulin delivery systems (Section 3.2).
Figure 3. Algorithmic security mechanism for GlucOS.
Figure 3 shows how our algorithmic security mechanism works. The whole process starts when a new CGM reading arrives. Once the BioKernel receives a new CGM reading, it collects a recent history of CGM readings and insulin dosing and sends these as inputs to both the reactive safe model and the predictive ML model. These models independently produce pump commands, called a temporary basal command. The temporary basal command sets the insulin delivery rate for the insulin pump for a 30-minute duration. The BioKernel will collect both temporary basal commands and look back at the differences between the reactive safe model and predictive ML model for the past three hours to see how much insulin delivery it should attribute to the predictive ML model. If the predictive ML model is still within its safe insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump. If it has exhausted its limits, we clamp it to fit within the insulin bounds we set relative to the reactive safe model.
To set bounds, we base our calculations on human physiology. People produce a background level of glucose for energy. People living with T1D need to inject insulin to facilitate the absorption of this background glucose.
Our method for setting bounds allows the predictive model to use up to three hours worth of background insulin in advance. If the prediction is correct and glucose levels rise (e.g., from eating), the proactive insulin will manage it. If the prediction is wrong and no extra glucose appears, the body’s ongoing glucose production will help use the extra insulin over the next three hours.
In the worst case, an incorrect prediction might lead to low glucose. However, with our mechanism in place, we limit the impact by our three-hour bound and offset by the body’s continuous glucose production.
3.2. Model principles
Our contribution is not in the specific algorithms we use, but rather in the insight around insulin delivery being equal for all correct algorithms and how we use this insight for security. In this section, we outline the algorithms we use in our current implementation and define the principles for how to apply security to any predictive insulin dosing algorithms.
Algorithmically, we use a PID controller (franklin2002feedback, ) from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model (mosquera2023enabling, ). Although we use this model in our current implementation, our system supports any predictive model. PID controllers are used in automated insulin delivery systems (weinzimer2008fully, ; kushner2018data, ) and research predictive ML models for insulin delivery have been well covered in the literature (Zhu2022, ; Jacobs, ; Zhang2021, ; Josep2020, ; RL, ; Tyler2022, ).
The design of the reactive safe model is crucial for algorithmic security, as it serves as the anchor for the security mechanism. The ideal reactive safe model should be safe, easy to understand, and have a formally verified implementation (Section 7).
To ensure the safety of this model, we base its calculations solely on facts observed automatically from the CGM and the insulin pump. Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much insulin they have injected but have not yet activated. From this calculation, we know how much insulin to inject or withhold.
These calculations are standard calculations that people who inject insulin manually use to determine dosing, making them easy for people who manage T1D or their medical care team to understand. The difference in our system is that we run these calculations automatically and every five minutes to adjust to the latest sensor readings. This basic formulation is a standard formulation used in most automated insulin delivery systems.
In contrast, predictive models have no restrictions since their outputs are constrained by the reactive safe model and our security mechanism. The ideal predictive model anticipates future metabolic states and adjusts insulin proactively to avoid glycemic imbalance.
In our evaluation, our predictive model outperforms the reactive safe model both in simulation (Section 9) and for the individual that uses our system for their real-world insulin dosing (Section 10), showing why it is important for automated insulin delivery systems to have a practical way to adopt ML.
4. Driver security: Withstanding malicious pump drivers
Pump drivers provide abstractions for insulin delivery applications to communicate with insulin pumps. Much like traditional device drivers, they have complex implementations since they help accomplish a wide range of tasks including pump setup/teardown, managing pump UX, handling insulin delivery/suspension, managing pump errors, accurately tracking insulin doses, and so on.
Given this complexity, there is a strong possibility for bugs or vulnerabilities to be present in existing pump drivers. These vulnerabilities can harm or even lead to the death of individuals since pump drivers control insulin pumps. For example, consider a scenario where the user decides to inject 3 units (3U) of insulin using the BioKernel. The BioKernel would relay this to the pump through abstractions provided by the pump driver. If the pump driver inadvertently or intentionally issues this command twice, 6U would be delivered, which could endanger the user by dosing too much insulin.
Our security goals for GlucOS’s driver security are to protect users from inadvertent bugs in pump drivers and safeguard against potentially malicious pump drivers. To achieve these goals, we introduce a shim between the pump driver and the pump, which validates all communications between them. We design our driver security mechanism to interpose on existing pump driver-pump communication instead of directly programming the pump. This design ensures that our trusted computing base remains small and simple by not including the complexity of pump driver implementations, thereby enabling formal verification. This design also provides flexibility with altering or introducing new functionality to pump drivers.
The most interesting aspect of our driver security mechanism was having to overrule our principle of not programming the pump in certain cases in the interest of protecting the individual. These cases involve commands pertaining to canceling ongoing insulin delivery. We could not directly adopt the design of other device driver security mechanisms (ref_validation, ; tang2010trust, ; expressos, ) for our driver security mechanism since these cases are specific to insulin pumps.
We focus our security efforts on pump drivers because they are complex (14.5k LoC), multi-threaded, event-driven, and use two-way communication between the pump and the driver. In contrast, CGM drivers are simple (1.9k LoC) and only read sensor values from the device. In our future work, we will formally verify our CGM driver using the same techniques we use for the insulin delivery path (Section 7).
4.1. Driver security mechanism design
Figure 4. We interpose on communications between the pump driver and the pump to handle untrusted pump drivers.
Figure 4 provides an overview of our driver security mechanism. The BioKernel registers high-level commands, like “deliver 2U of insulin” with our driver security mechanism, and sends them to the pump driver. The pump driver converts these commands into I/O messages to program the pump. The driver security mechanism then checks to make sure that these I/O messages are consistent with the previously registered high-level pump commands before forwarding them to the pump hardware. We only validate commands for insulin delivery since they pose the biggest threats to the user. On detecting added, modified or dropped commands from the pump driver, the driver security mechanism rejects further commands and transitions the system to a manual mode where the user takes control.
Dropped commands pertaining to canceling ongoing insulin delivery requires separate consideration since excess insulin is dangerous and cannot be withdrawn once it has been delivered. Users will cancel insulin delivery if they accidentally issued an insulin delivery command or incorrectly set the insulin dosage. In cases where a malicious driver drops a cancel command, the driver security mechanism directly issues I/O messages to the pump to cancel insulin delivery before transitioning the pump to a manual mode. We make the tradeoff of increasing the size of our trusted computing base to handle these cases in the interest of protecting the user. Fortunately, there are only two commands to cancel insulin delivery in our current implementation, so the increase to our trusted computing base is small.
5. End-to-end security: Biological invariant
When implementing security systems for humans, even the perfect algorithmic and driver security mechanisms are unable to handle all situations due to changes in human physiology. A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state. If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong. These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin (borghouts2000exercise, ), or if the pump insertion site hits a vein and delivers insulin directly into the blood stream (steyn2023accelerated, ). Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user. We call this safety check that accounts for changing human physiology a biological invariant.
In our implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs. If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be. We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detection of significant glycemic deviations above this level (30mg_1, ; 30mg_2, ).
We take a three-tiered approach to dealing with deviations in our glucose change calculations. First, for small differences of less than 30 mg/dl, we accumulate errors and adjust the target glucose level to compensate, clamping the adjustment to 10 mg/dl. For example, if the individual’s insulin sensitivity is elevated and their target glucose level is 90 mg/dl, GlucOS will update their target to 100 mg/dl dynamically to adjust. Second, when the differences are 30 mg/dl or greater, we shut off insulin delivery. Our rationale is that the individual is more sensitive to insulin than our models believe they should be, thus we want to avoid overdosing. If their difference goes back to within the expected range, normal execution resumes. Third, if they remain at 30 mg/dl or more for more than two hours, we alert the individual and transition the system into a manual mode that requires the user to dose insulin manually.
6. Last line of defense: Humans
In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system. While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance. We discuss our most interesting findings around preferences and the architectural iterations we make to our design based on our real-world experiences.
6.1. Predictive alerts for humans
When an individual’s glucose levels are not within the range of healthy values, they need to intervene to bring them back in range. Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered. To recover from hypoglycemia, people need to eat sugar. Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
Figure 5. GlucOS predicts hypoglycemia and notifies the user before it happens so they can proactively prevent it.
We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 5), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack. Alerts for hypoglycemia are not new. All CGM software comes with non-predictive hypoglycemia alerts by default (dexcom_alerts, ) and glucose prediction algorithms have been well covered in the literature (Josep2020, ; Zhang2021, ; arora2023multivariate, ). Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
In a user study (Section 10.4), we reduced the amount of time six people spent in hypoglycemia as a result of using GlucOS, despite them already having CGM alerts set up prior to our study. These results demonstrate a reduction in hypoglycemia for the non-attack case, but the most dangerous successful attacks will result in hypoglycemia. Thus, these results suggest that predictive alerting can be an effective security measure.
Using a simple and deterministic algorithm is important since it runs as a part of our trusted computing base. Algorithmically, we predict hypoglycemia by doing a least squares fit for the last 20 minutes of CGM readings and extrapolate forward 15 minutes. We trigger alerts if our prediction shows hypoglycemia is likely in the next 15 minutes.
The most interesting finding from our study was that personalization was important. We provided people with the ability to set the prediction threshold and specify how often they wanted the alert to repeat while they remained out-of-range. Five of the six participants customized their alerts and the participant noted that the default values we picked coincided with their alerting preferences. In our first implementation we did not provide the ability to customize their security settings and we tried to make the usability vs security tradeoff algorithmically and automatically. But in the end, simplifying the system and allowing users to personalize their security policy was important for our participants.
Our finding has the potential to have broader implications for security in healthcare systems. Often, security systems favor uniformity and consistency (DHILLON2016656, ; FURNELL20165, ), but when managing people’s health we found that allowing people to balance alert fatigue and security increased engagement and potentially increased protection, based on each individual’s needs.
6.2. Architectural implications
One interesting design iteration we made was moving the alerting logic into the BioKernel itself. In our first design, consistent with our separation principles, the alerting logic resided in a separate app. This alerting app learned the state of the system using our cloud-based event logging system. However, one of our participants went on a canoeing trip and was disconnected from the internet, so they could not get alerts. Ideally iOS would provide inter-process communication mechanisms to pass data from the BioKernel directly to the alerting app, but iOS does not have the appropriate APIs for this style of communication. Thus, we moved the alerting logic into the BioKernel, violating our principle of separation, but enabling our users to get alerts even when they are disconnected from the internet.
Moving the alerting logic into the BioKernel did have security implications. It adds complexity to our trusted computing base, but we carefully manage this tradeoff. We minimize the added complexity by using traditional, well-understood algorithms for prediction rather than more complex ML models and reuse existing functionality within the BioKernel when possible to keep the implementation simple. This approach allowed us to maintain a high level of confidence in the security and correctness of our alerting system while still providing the necessary functionality.
7. Formal verification
To provide additional assurance that our implementation is correct, we port our Swift code to Dafny and use Hoare logic. As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example (chong2020code, ). Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent. We port updated code back to Swift manually because Dafny does not include code generation for Swift.
These efforts paid off: we found and fixed nine bugs in our original implementation; in addition to proving various invariants in the code for our insulin delivery path and security mechanisms. However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms.
Our implementation operates on CGM readings, but in our application of formal methods, our goal is to reason about the individuals actual glucose level, taking into account sensor errors. This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters that are part of the software implementation. We prove that the biological invariant holds, given error bounds set by the CGM manufacturer (garg2022accuracy, ).
To reason about actual glucose levels formally, we use a ghost variable in Dafny. Ghost variables are variables that are only available to the prover, these do not exist in the real software system. By defining the actual glucose levels using a ghost variable, we can include it in preconditions and postconditions to prove properties about our software relative to actual glucose levels.
We believe that having a parameter for the real biological state of the system is an important modeling decision; and it diverges, for example, from previous work that only models the biological data in simulation (zhou2021data, ; zhou2023knowsafe, ). Traditional software verification techniques like model checking (cps_1, ) and deductive verification (cps_3, ) have been explored in the past for hybrid system verification (cps_2, ) that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process. Also, researchers have used real-world invariants for security in the context of autonomous vehicles (choi2018detecting, ), but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
8. Implementation
In our implementation, the BioKernel is an iOS app that consists of 5.1k lines of code in our trusted computing base. It communicates with a Dexcom G6 or G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy, and runs the core closed-loop algorithm. For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state. Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system (Figure 1(c)) that enables separate apps to extend functionality while remaining isolated from the BioKernel.
9. Evaluation
In this section, we evaluate GlucOS’s algorithmic and driver security mechanisms. In the next section (Section 10), we evaluate all components of GlucOS in the real world.
9.1. How effective is GlucOS’s algorithmic security mechanism?
We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator (simglucose, ). GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D. We consider extreme cases of attack in our evaluation. While subtle attacks are also possible, we focus on extreme attacks since they have the most adverse consequences.
Concreteny, we first run simulations with a predictive ML model inspired from prior research (mosquera2023enabling, ), without incorporating our security mechanism to set a baseline of an effctive algorithm in managing T1D. Then, in the same simulation scenarios, we introduce an order of magnitude increase to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms. We then repeat these experiments with our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
The American Diabetes Association (ADA) recommends that individuals with T1D spend at least 70% of their time in the range (TIR) of 70-180 mg/dl to remain healthy (nathan2014diabetes, ; endocrine_tir, ; journal_glycemic_targets, ). Glucose levels less than 70 mg/dl (hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels greater than 180 mg/dl (hyperglycemia) increase risks for long-term vascular damage.
9.1.1. Without algorithmic security mechanism
Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training. Figure 6 shows the average proportion of time spent in the safe target range (TIR) by virtual humans from different age groups when employing the predictive ML model. We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model. These results provide a baseline for an algorithm that is effective in managing T1D.
Reactive TIR with TIR with
safe predictive clamped
Age group model ML model ML model
Adults 88.25% 93.49% 92.88%
Adolescents 81.98% 85.23% 82.38%
Children 77.31% 86.6% 85.42%
Figure 6. Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms. In this section, we discuss an experiment with a malicious algorithm that intentionally doses too much insulin. Concretely, our malicious algorithm doses ten times the amount of insulin computed by the previously described predictive ML model. We ran this algorithm within the simulator on the same virtual human on the same scenarios as the previous experiment.
Figure 7 shows the CGM readings on three randomly chosen virtual humans on a particular scenario when we dose more. The CGM readings of all 3 individuals fully go below the hypogycemic bound, which would have killed them in the real-world.
Figure 7. Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin. The CGM readings staying below the hypoglycemic bound would have likely been fatal to all three individuals.
Across all virtual humans, we see an average drop in TIR from 88.44% to 0% when dosing too much insulin. Correspondingly, we observe an average increase in the time spent in hypoglycemia from 0% to 100%. We also observed similar results with experiments where we intentionally underdosed insulin. In these experiments, all virtual humans spend most of their time above the hyperglycemic bound which wouldn’t have been fatal but would have led to severe health consequences in the real world. These results motivate the need for our algorithmic security mechanism.
9.1.2. With algorithmic security mechanism
Impact on incorrect dosing. We repeat experiments from the previous section with the same scenarios. We run the same malicous model to inject too much insulin, but also run our reactive safe model (Section 3) to bound insulin injection.
Figure 8 show CGM readings on the same three virtual humans on the same scenarios as the experiment in the previous section. We see that all of them spend over 70% of their time in range, showing that GlucOS protects individuals even in presence of malicious overdosing. On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 78.22% when employing safety while overdosing. We also report similar results on experiments with a malicious algorithm that underdoses insulin.
While the times spent in range are less than those of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing active attacks.
Figure 8. Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin. On average, CGM readings stay in range over 70% of the time.
Impact on effective benign algorithms. We evaluate the impact of GlucOS’s algorithmic security mechanism on benign algorithms that are effective in managing T1D such as the baseline predictive ML model. We repeat the experiment of running the predictive ML model on all virtual humans in the same scenarios with the same parameters for the clamps and reactive model used to evaluate incorrect dosing.
Comparing data in Figure 6 we see that there is only a slight drop in the average proportion of time spent in range by virtual humans across all age groups when employing GlucOS’s algorithmic security mechanism. When employing the reactive safe model as is to calculate insulin doses, we report an average TIR of 82.48% which is lower than the average TIR of 86.84% observed when running the predictive ML model with GlucOS’s security mechanism. This shows that GlucOS’s algorithmic security guarantees do not come at the cost of effective T1D management from predictive ML.
Figure 9. We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands. Our driver security mechanism detected all manipulations.
9.2. Does GlucOS ensure pump operation while overcoming malicious pump drivers?
Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver.
Impact on malicious pump drivers. We attached an Omnipod DASH insulin pump to a stuffed animal (Figure 9). We then modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands. We detected all manipulated commands thereby showing that our driver security mechanism has 0 false negatives in detecting them.
Impact on benign pump drivers. For one week, we logged messages passed between the BioKernel, the unmodified OmniBLE pump driver and the pump on an individual running GlucOS (Section 10). We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanism does not incur any false positives.
These results show that GlucOS can ensure pump operation while withstanding attacks from malicious drivers.
10. Using GlucOS in the real world
In this section, we report on our experiences working with seven individuals using GlucOS to help manage their T1D. Six people are students who only used the predictive alerting feature of the BioKernel (Section 6) for two months, from June 2024 - August 2024. One individual, who we will call Bob, has been using the full GlucOS system since November 2023 for insulin delivery and predictive alerting.
10.1. Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe. Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They designed the safety protocol, which defines the specific criteria we use to stop the study, if needed. The Appendix will include a more detailed discussion of our ethics and safety protocols.
10.2. Is an ML-based closed-loop system safe and effective in practice?
Bob has been using GlucOS in a closed loop since November 23rd, 2023. From November 23 to January 28th, 2024 we used our reactive safe model to make control decisions. Starting on January 28th Bob started to use our predictive ML model for periods of time to get used to the new system, and then starting on January 31st, 2024 starting using our predictive ML model exclusively. We report on one week’s worth of data from the predictive ML model starting on January 31st ending on February 6th, 2024.
Overall, Bob had the best results when using ML. His time spent in range increased from 96.3% to 97.2%, and notably he spent no time in hypoglycemia (below 70 mg/dl), down from 0.67% when not using ML.
Next, we measure how often the predictive ML model was responsible for programming the pump vs our reactive safe model. For the one week when Bob was using the predictive ML model, 33.0% of the time both the predictive ML model and reactive safe model issued the same commands, 39.9% of the time the BioKernel used the predictive ML model’s commands, and 27.1% of the time the BioKernel used the reactive safe model’s commands. We believe that this distribution represents a suitable balance for providing the ML with enough flexibility to improve outcomes, while having tight enough bounds to limit potential damage from bad or malicious ML predictions.
10.3. Do biological invariant violations trigger in practice?
To determine if biological invariant violations trigger in practice, we review 90 days of data for Bob, from May 29th, 2024 to August 27th, 2024. During this period of time, biological invariant violations occurred 1.6k times, or 9.0% of the time. Of these violations there were 49 sequences of 30 minutes or longer with the longest sequence being 65 minutes.
These results suggest that although natural variations in glucose metabolism do lead to biological invariant violations, none of them were long enough to trigger our most severe action of transitioning to manual mode after two hours.
We reviewed one weeks worth of violations with Bob manually to determine the cause. During the most recent week of data, Bob had five sequences of 30 minutes or longer and all were caused by exercise.
10.4. Are predictive alerts useful for security?
To measure predictive alerts as a security mechanism (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts. Our primary metric is time spent below 70 mg/dl, where a decrease in the amount of time spent in hypoglycemia suggests that predictive alerts are effective to empower people to mitigate security attacks.
For the two month study, our participants decreased their average time in hypoglycemia from 2.7% down to 1.6%, and all individuals experienced a decrease in the amount of time spent in hypoglycemia. For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% (brown2019six, ), showing that our 1.1% decrease is substantial.
Overall, our results suggest that predictive alerts can be a useful last line of defense for trustworthy automated insulin delivery systems.
10.5. How does the BioKernel perform?
We look at two aspects of system performance when evaluating our real-world use of GlucOS. First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end. We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire eleven month period.
On average, the BioKernel spends 4.3s running the closed loop algorithm, where almost half of that time is dominated by CGM and pump communications. Of the components within the closed-loop algorithm, the ML predictions take 2.2s on average and the safety logic takes 19ms (0.019s). Since the system runs in the background, we believe that iOS runs the device with reduced computational capacity to save power. But overall, 4.3s of wall-clock time for every five minutes our closed-loop algorithm runs with only a tiny fraction of that coming from our safety logic represents practical system performance.
For successful closed-loop end-to-end execution, over the evaluation period the closed-loop algorithm runs successfully 96.5% of the time, with 2.1% of the runs failing due to pump communication, which is our largest class of failure.
11. Related work
In addition to the work already discussed in this paper, there are several other related works.
The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount. Current research on security for these systems focuses primarily on evaluating algorithms using simulations (zhou2021data, ; zhou2023knowsafe, ). We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints. Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system. Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components. Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable. These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems. To address these challenges, GlucOS moves beyond simulation-based designs and tackles the challenges of real-world systems.
Previous research has looked at the security of implanted medical devices in general (halperin2008security, ; burleson2012design, ; rushanan2014sok, ), in addition to looking at insulin pumps in particular (li2011hijacking, ; paul2011review, ), with more recent work looking at providing improved security (marin2016feasibility, ; ahmad2018securing, ). Also, recent work has looked at applying formal methods to insulin pumps for high assurance (panda2021secure, ). These works focus on devices and their communication channels. In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
12. Conclusion
The design of GlucOS, while grounded in established OS and security principles, incorporates several non-obvious elements that emerged from our deep engagement with the problem domain and real-world testing.
First, our approach to algorithmic security used classic security techniques, like reference monitors and the Simplex architecture. However, it was our novel insight around all correct algorithms dosing the same amount of insulin that led us to these classic techniques. Plus, from our exploration, the most interesting part was not the ML, it was around the principles of reactive safe models and in particular our principle that these safety-critical algorithms should be reactive.
Second, our pump driver security went beyond traditional specification and checking. We needed to actively program the pump to handle dropped cancel commands, which was an unexpected but crucial insight.
Third, the biological invariant emerged as a key check in our system, addressing a blind spot in typical simulations that ignore changing human dynamics. This forced us to develop novel formal methods to handle scenarios where our glucose metabolism models could be drastically wrong.
Fourth, our approach to human interaction in the system was nuanced. We found that simply alerting humans or suspending insulin delivery were not always appropriate. For security alerts, personalization and customization were important. Stopping insulin delivery could be dangerous if the user has high glucose and insufficient insulin on board, which required more context-aware responses.
These non-obvious design decisions underscore the complexity of creating a truly trustworthy automated insulin delivery system. They reflect the value of combining theoretical security principles with practical, real-world testing and domain-specific knowledge. This interdisciplinary research drew on deep expertise from computer security, medicine, and ethics to create a practical system.
Our exploration is an extreme example of personalized, distributed, and automated healthcare using trustworthy computer systems.